# Benchmarking Multimodal LLMs on Recognition and Understanding over Chemical Tables

### v1

@@ Abstract
Chemical tables encode complex experimental knowledge through symbolic expressions, structured variables, and embedded molecular graphics. Existing benchmarks largely overlook this multimodal and domain-specific complexity, limiting the ability of multimodal large language models to support scientific understanding in chemistry. In this work, we introduce ChemTable, a large-scale benchmark of real-world chemical tables curated from the experimental sections of literature. ChemTable includes expert-annotated cell polygons, logical layouts, and domain-specific labels—including reagents, catalysts, yields, and graphical components—and supports two core tasks: (1) Table Recognition, covering structure parsing and content extraction; and (2) Table Understanding, encompassing both descriptive and reasoning-oriented question answering grounded in table structure and domain semantics. We evaluated a range of representative multimodal models, including both open-source and closed-source models, on ChemTable and reported a series of findings with practical and conceptual insights. Although models show reasonable performance on basic layout parsing, they exhibit substantial limitations on both descriptive and inferential QA tasks compared to human performance, and we observe significant performance gaps between open-source and closed-source models across multiple dimensions. These results underscore the challenges of chemistry-aware table understanding and position ChemTable as a rigorous and realistic benchmark for advancing scientific reasoning  [^1] https://github.com/lqzxt/ChemTable.
@@ 1 Introduction
Recent advances in multimodal large language models (MLLMs) have created new opportunities for mining expert knowledge from scientific literature and are increasingly viewed as catalysts for AI-driven scientific discovery [1]. A growing wave of scientific accelerators—such as MLLMs-based OCR, ChatPaper [2], and ChatPDF [3]—demonstrates the potential of MLLMs to streamline literature parsing, summarization, and interactive reading. From a modeling perspective, MLLMs exhibit strong capabilities in semantic understanding and reasoning, making them well-suited for processing the rich multimodal content of scientific documents [4, 5]. Benchmarks like ChartQA [6] and ChartX [7] have begun exploring visual reasoning over scientific figures. Yet, scientific literature remains one of the most semantically dense and domain-specialized corpora, and understanding it serves as both a valuable application and a rigorous testbed for evaluating the limits of MLLMs [8].
@@
Figure 1: Multimodal Complexity of Chemical Tables and Challenges for MLLMs.
Among scientific tables, those found in chemistry pose especially complex challenges that go far beyond standard layout parsing [9]. A typical chemical table encodes rich experimental workflows through dense symbolic notation (e.g., reagents, ligands), domain-specific abbreviations (e.g., “BINAP”, “TFA”), and embedded visual elements such as molecular structures or reaction schemes [10, 11]. As illustrated in Figure 1, a single row often represents a multi-variable configuration—catalyst, ligand, solvent, additive, with quantitative outcomes like yield, selectivity. These tables also rely on implicit conventions (e.g., ratio formats like “$>$19/1”) and footnoted exceptions, making their semantics both subtle and compact. Interpreting such tables requires aligning symbolic, numeric, and visual information in a domain-aware manner, which presents significant challenges for general-purpose MLLMs not trained on scientific representations [8, 9].
To systematically address these challenges, we introduce ChemTable, a high-quality benchmark designed for recognition and understanding over real-world chemical tables. The dataset comprises over 1,300 tables curated from peer-reviewed chemistry literature, spanning diverse reaction types, experimental conditions, and reporting formats. ChemTable supports two core tasks: table Recognition[12]—including structure reconstruction and content extraction—and table Understanding [13, 14], formulated through more than 9,000 QA instances across two categories: (1) descriptive questions, which evaluate a model’s ability to extract key facts; and (2) reasoning questions, which require comparison, attribution, and domain-grounded inference.
To enable scalable and consistent evaluation, all answers follow a short-form format suitable for MLLMs-based automatic grading. We benchmark seven widely-adopted MLLMs, including representative models from both open- and closed-source sources, and observe significant performance gaps across all tasks. Compared to human and expert performance, current MLLMs lag behind—particularly in interpreting symbolic and embedded graphical elements. Our analysis further reveals key insights, including the symbolic understanding gap, the limited transferability of MLLMs to scientific domains, and substantial variation across model families. We release ChemTable, along with evaluation tools, to support future research in multimodal scientific understanding.
@@ 2 Related Work
Existing table recognition benchmarks mainly focus on two tasks: Table Structure Recognition (TSR) and Table Recognition (TR). The widely used ICDAR-2013 [15] dataset includes 257 manually annotated tables for TSR. Despite its small size, it remains a key benchmark due to its annotation quality. Since 2019, large-scale TR datasets [16, 17, 18, 19] have emerged. However, their annotations are programmatically generated and provide only table structures without cell content, limiting their utility in deeper understanding tasks. To address this, FinTabNet [20] and PubTabNet [21] incorporate logical cell locations, enhancing TSR modeling. SciTSR [22] focuses on scientific documents, offering detailed cell content annotations, thus better supporting academic table recognition.
@@
Several datasets have been proposed for table understanding tasks, including WikiTQ [23] and WikiSQL [24]. These datasets consist of numerous tables extracted from Wikipedia and serve as critical early benchmarks in table understanding research. In recent years, advancements in natural language processing have extended beyond traditional homogeneous tables to incorporate additional modalities. For instance, HybridQA [25] and MMTab [26] introduce multi-modal or semi-structured sources for more complex reasoning. Moreover, domain-specific benchmarks such as FinQA [27] (financial) and SciTab [28] (scientific) address unique challenges by integrating structured tables with related textual content to support complex reasoning tasks.
@@
WikiTQ [23]
WikiSQL [24]
HybridQA [25]
FinQA [27]
SciTab [28]
MMTab [26]
@@
The ChemTable dataset is different from existing benchmarks in several key ways. It is a comprehensive, open-ended dataset focused on table recognition and understanding in chemistry, created to support scientific question-answering systems. Unlike other datasets, its questions combine visual elements, table data, and chemistry knowledge, making it necessary for models to recognize table content and structure, then reason across that information. This setup reflects real tasks that researchers face when drawing conclusions from data. ChemTable also helps evaluate how well multimodal language models can reason, showing their strengths and weaknesses.
@@ 3.1.1 Table Collection
Journal Selection. The journal selection criteria prioritised top-tier chemistry journals (e.g., ACS Catalysis, JACS, Chem, Angewandte Chemie Int. Ed., and Science) to ensure data credibility and disciplinary relevance. Publications from the past five years (2015–2025) were prioritised to balance cutting-edge insights with foundational studies, while DOI links were systematically documented to ensure traceability and facilitate future validation or expansion of the dataset.
@@ 3.1.2 Table Annotation
For table annotation, we formed a dedicated team to identify table titles, annotations, main content, and additional captions [29]. Each element was annotated with pixel coordinates and OCR-verified text. We also encoded the logical structure of tables, such as row-column relationships. To preserve visual meaning, we carefully recorded stylistic features like boldface, italics, and color. For table-based question answering annotation, descriptive questions are directly derived from the detailed annotations of table elements created in the previous step. Simple reasoning questions are generated using GPT-4.1 [30] and then filtered based on difficulty. For more complex reasoning questions and visually descriptive tasks, we manually annotated 2,122 questions with the help of experienced graduate students. These questions focused on specific topics, such as reaction conditions and yields, while allowing a variety of question styles. Annotators were also encouraged to include unanswerable questions caused by missing data, vague references, or incomplete formatting. To ensure accuracy, all questions were verified by human review and model checks. Details can be found in the Appendix.
@@ 3.2 Table Recognition
Table recognition is a fundamental step in document understanding, which aims to turn table images into structured data. This task is more challenging in chemistry because of complex layouts and specialized symbols [31].
@@ 3.2.2 Evaluation Protocols
Molecular Recognition: Chemical tables often include molecular structures represented as images, either embedded within cells or positioned externally. This task aims to evaluate a model’s ability to recognise and interpret such molecular graphics. The objective is to extract the corresponding SMILES (Simplified Molecular Input Line Entry System) string from a molecular diagram. This task presents unique challenges, such as fine-grained visual understanding and domain-specific symbol interpretation, which are not typically encountered in general-domain table recognition [32, 33].
Table 2: ChemTable Dataset Statistics. Unique Tokens and Question & Answer Lengths are Calculated Based on the Qwen-7B Tokenizer.
@@ 3.3.1 Task Definition
Descriptive questions aim to provide a general overview of the basic information presented in the table shown in the image. These questions include: 1) describing the main body of the table, such as its dimensions; 2) describing basic metadata of the table, including titles and notes; 3) describing domain-specific elements, such as reaction conditions in chemical reaction formulas or SMILES notation in molecular graphs; and 4) identifying certain visual features in the table, such as rows highlighted with special colours. Chemical tables often contain images, such as molecular structures, instead of plain text, which makes them harder to read and analyse. For example, an image in a table row can affect spacing and make it harder to recognize content. Sometimes, only certain parts of a molecule, like a red-highlighted –OH group—are shown, which adds to the difficulty for models trying to understand the table.
@@ 3.3.3 Reasoning Questions
Reasoning questions are designed to further analyze and infer information from the data presented in the table within the image. They require a comprehensive understanding to make informed judgments. These questions include: 1) Numerical and statistical reasoning, 2) Trend and change analysis, 3) Multi-hop logical reasoning, and 4) Domain-specific reasoning. Specific details are provided in the Appendix. Since the data in chemical tables is often tightly linked to specific experimental conditions, molecular structures, and graphical annotations, reasoning questions not only assess the ability to comprehend explicit information but also place greater demands on domain knowledge, the integration of information across rows and columns, and the ability to reason using visual symbols.
@@ 3.3.4 Data Filtering
Diversity. We identified questions with overly repetitive structures and semantically similar phrasing to encourage a broad range of question formulations. These were rewritten using GPT-4.1 with prompt templates provided in the Appendix. We selected algorithms that maximised the semantic distance from the original question, enhancing the linguistic and structural diversity of the dataset.
Difficulty. We implemented a filtering strategy to ensure the dataset poses a meaningful challenge. We first conducted a single-pass QA evaluation using the Qwen-2.5-7B model [34]. For each question, we recorded whether the model was able to produce the correct answer on the first attempt. Questions that were answered correctly in one try were deemed too simple. To filter out these low-difficulty samples, we randomly discarded some of them. This approach allowed us to enrich the dataset with more challenging examples that were better suited for evaluating advanced reasoning capabilities.
@@ 4.1 Experimental Setup
To evaluate the performance of the table recognition task, we adopt the improved similarity metric based on tree edit distance (TEDS) [21] along with the TEDS-structure indicator. Specifically, the table content may contain molecular graphs in chemical scenarios. Interpreting these molecular graphs using the simplified molecular input line entry system(SMILES) can result in structural isomorphism, where different representations or atom orders may correspond to the same molecule, making it inappropriate for TEDS to utilise normalised edit distance as a measure of cell content recognition accuracy. Therefore, for cells containing chemical molecular graphs, we replace the normalised edit distance with the Tanimoto coefficient [35] to accurately assess the performance of table recognition. We use accuracy (ACC) as the evaluation metric for fine-grained retrieval experiments.
@@
We evaluate a diverse set of MLLMs. For open-source models, we test InternVL3-78B [36], Llama-3.2-90B [37], and Qwen2.5-VL-72B [38]. We also evaluate the following proprietary models: Gemini-2.5-Flash [39], GPT-4.1, GPT-4.1-mini [30], and Claude-3-7-Sonnet [40].
@@ 4.2 Experimental Results
We evaluated MLLMs on molecular formula recognition from real-world academic papers and synthetic diagrams. As shown in Figure 3, MLLMs can identify and convert molecular structures into chemical formulas, but their performance on real-world diagrams is lower than on synthetic ones, highlighting the impact of data diversity and quality on model robustness. Although MLLMs possess some chemical domain knowledge, their accuracy and reliability remain significantly inferior to specialized models (DECIMER[41]). This gap underscores the need for further improvements to enable MLLMs to effectively handle domain-specific tasks in advanced scientific applications.
@@ 5.1 Experimental Setup
For tasks requiring descriptive answers, we use edit distance [42] as the primary metric to assess recognition accuracy. For more open-ended question answering (QA) tasks, we adopt a binary evaluation strategy powered by GPT-4.1-nano [30], which classifies each response as either correct or incorrect [43, 44, 45],. The overall performance is then quantified by computing the accuracy (ACC) based on these binary classifications.
To ensure the reliability of this automated evaluation process, we randomly sampled 20% of the QA instances for manual verification. Human annotators reviewed the model outputs against reference answers and provided independent judgments. We then calculated the agreement rate between human evaluations and the binary assessments produced by GPT-4.1-nano . Implementation details are provided in the Appendix.
@@
We evaluate a comprehensive set of general-purpose multimodal large language models (MLLMs), including both open-source and proprietary systems. The open-source models evaluated are InternVL3-78B (InternVL) [36], Llama-3.2-90B (Llama) [37], and Qwen2.5-VL-72B (Qwen-VL) [38]. Among the proprietary models, we assess Gemini 2.5 Flash (Gemini) [39], GPT-4.1, GPT-4.1 mini (GPT-mini) [30], and Claude-3.5-Sonnet (Claude) [40]. To provide an upper-bound reference, we also include a human performance baseline (Human), which consists of answers manually provided by annotators on a representative subset of the evaluation data.
@@
Question Type GPT-mini GPT-4.1 Claude Gemini Qwen-VL InternVL Llama Human
@@
Element Description Table Dimensions 68.31 73.07 75.39 73.10 71.42 70.50 70.27 -
Title Description 84.79 87.31 81.03 81.64 83.18 84.35 85.50 -
Annotation Description 81.86 92.94 68.93 81.12 90.23 73.27 81.39 -
Molecular Recognition 37.62 42.49 52.71 28.47 14.14 11.63 18.50 -
Visual Description Bold Description 35.88 44.27 50.38 41.22 38.93 32.82 52.73 98.99
Color Description 54.92 56.48 50.38 41.22 58.55 49.74 58.03 97.73
@@
Domain- Specific QA Benzene Rings Count 49.51 52.31 62.83 75.32 59.61 47.66 21.97 94.98
Yield and Conditions 89.97 89.14 89.42 93.87 85.24 74.93 82.13 93.61
Function Based 20.78 37.30 62.06 71.70 35.37 30.23 25.83 89.27
Numerical Statistics Value Comparison 78.40 91.80 93.60 98.40 94.40 67.14 81.45 100.00
Find Min/Max 79.44 85.85 94.39 83.18 94.39 60.94 60.95 100.00
Calculate Sum 32.63 58.33 53.68 64.21 32.63 24.07 34.12 100.00
Calculate Average 22.52 44.87 66.67 72.97 26.13 24.19 33.33 98.20
Logical Reasoning Trend Analysis 75.13 81.87 83.94 88.60 74.61 76.34 55.96 98.45
Multi-hop Retrieval 83.55 84.87 88.16 87.94 82.89 80.20 81.48 98.67
@@
We compare our method against a set of representative MLLMs on table understanding task in Table 4. Below are our main findings:
(a) General reasoning strength, numerical weakness. MLLMs achieve high accuracy (ACC > 0.8) on tasks such as trend analysis, value comparison, and finding min/max values. Specifically, Gemini scores 98.40 on value comparison and 88.60 on trend analysis. This indicates strong capability in capturing basic quantitative and logical patterns from tables. However, performance drops significantly on arithmetic-heavy tasks like calculating the sum or average, where Gemini only reaches 64.21 and 72.97, respectively, highlighting limitations in calculation-intensive reasoning.
(b) Descriptive tasks outperform domain-specific chemistry QA. MLLMs show strong performance in general descriptive tasks. GPT-4.1 achieves 92.94 on annotation description. In contrast, accuracy plummets on domain-specific chemistry questions. Gemini scores only 75.32 on benzene ring counting and 71.70 on function-based QA. This reveals limitation of current models in handling specialized chemical content, especially when integrating domain knowledge with visual input.
(c) Visual style interpretation remains challenging. Tasks that rely on recognizing formatting cues — such as bold or color highlights — yield consistently low accuracy (e.g., Llama scores 52.73 on bold description and 58.03 on color description). Even top-performing models struggle to interpret these stylistic features embedded within structured data, suggesting that visual structure understanding is still an underdeveloped capability in current MLLMs.
(d) Closed-source models dominate complex and domain tasks. Closed-source models like Gemini and GPT-4.1 consistently achieve top results across most categories — Gemini leads in value comparison (98.40), trend analysis (88.60), and several domain-specific tasks. In contrast, open-source models like InternVL and Llama lag behind, particularly in complex reasoning and chemical QA, indicating that current open models are not yet competitive in specialized, vision-language reasoning domains without further fine-tuning or domain adaptation. Additionally, despite impressive performance, these models still fall slightly short of human capabilities in complex tasks.

### v2

@@ Abstract
With the widespread application of multimodal large language models in scientific intelligence, there is an urgent need for more challenging evaluation benchmarks to assess their ability to understand complex scientific data. Scientific tables, as core carriers of knowledge representation, combine text, symbols, and graphics, forming a typical multimodal reasoning scenario. However, existing benchmarks are mostly focused on general domains, failing to reflect the unique structural complexity and domain-specific semantics inherent in scientific research. Chemical tables are particularly representative: they intertwine structured variables such as reagents, conditions, and yields with visual symbols like molecular structures and chemical formulas, posing significant challenges to models in cross-modal alignment and semantic parsing. To address this, we propose ChemTable—a large-scale benchmark of chemical tables constructed from real-world literature, containing expert-annotated cell layouts, logical structures, and domain-specific labels. It supports two core tasks: (1) table recognition (structure and content extraction); and (2) table understanding (descriptive and reasoning-based question answering). Evaluation on ChemTable shows that while mainstream multimodal models perform reasonably well in layout parsing, they still face significant limitations when handling critical elements such as molecular structures and symbolic conventions. Closed-source models lead overall but still fall short of human-level performance. This work provides a realistic testing platform for evaluating scientific multimodal understanding, revealing the current bottlenecks in domain-specific reasoning and advancing the development of intelligent systems for scientific research. [^1] https://github.com/lqzxt/ChemTable
@@ 1 Introduction
Recent advances in multimodal large language models (MLLMs) have created new opportunities for mining expert knowledge from scientific literature and are increasingly viewed as catalysts for AI-driven scientific discovery (zhang2024comprehensive). A growing wave of scientific accelerators—such as MLLMs-based OCR, ChatPaper (dean2023chatpapers), and ChatPDF (panda2023enhancing)—demonstrates the potential of MLLMs to literature parsing, summarization, and interactive reading. From a modeling perspective, MLLMs exhibit strong capabilities in semantic understanding and reasoning, making them well-suited for processing the rich multimodal content of scientific documents. Benchmarks like ChartQA (masry2022chartqa) and ChartX (xia2024chartx) have begun exploring visual reasoning over scientific figures. Yet, scientific literature remains one of the most semantically dense and domain-specialized corpora, and understanding it serves as both a valuable application and a rigorous testbed for evaluating the limits of MLLMs (li2024scilitllm). In particular, while MLLMs excel at general-purpose visual tasks, they continue to struggle with domain-specific multimodal reasoning, precisely the capability that underpins AI-assisted scientific discovery.
@@
Figure 1: Illustrative examples from the ChemTable dataset, showcasing the diverse and multimodal challenges for Multimodal Large Language Models.
Among scientific tables, those found in chemistry pose especially complex challenges that go far beyond standard layout parsing (abdelmagid2014survey). A typical chemical table encodes rich experimental workflows through dense symbolic notation (e.g., reagents, ligands), domain-specific abbreviations (e.g., “BINAP”, “TFA”), and embedded visual elements such as molecular structures or reaction schemes (chemicalentity; smicrm). As illustrated in Figure 1, a single row often represents a multi-variable configuration—catalyst, ligand, solvent, additive, with quantitative outcomes like yield, selectivity. These tables also rely on implicit conventions (e.g., ratio formats like “>>19/1”) and footnoted exceptions, making their semantics both subtle and compact. Interpreting such tables requires aligning symbolic, numeric, and visual information in a domain-aware manner, which presents significant challenges for general-purpose MLLMs not trained on scientific representations (li2024scilitllm; abdelmagid2014survey).
To systematically address these challenges, we introduce ChemTable, a high-quality benchmark designed for recognition and understanding in chemical tables. The dataset comprises over 1,300 tables curated from peer-reviewed chemistry literature, spanning diverse reaction types, experimental conditions, and reporting formats. ChemTable supports two core tasks: table Recognition(survey2025)—including table structure reconstruction and content extraction—and table Understanding (surveytabular2024; surveyour), formulated through more than 9,000 QA instances across two categories: (1) descriptive questions, which evaluate a model’s ability to extract key facts; and (2) reasoning questions, which require comparison, attribution, and domain-grounded inference.
To enable scalable and consistent evaluation, all answers follow a short-form format suitable for MLLMs-based automatic grading. We benchmark seven MLLMs for table recognition and ten MLLMs for table understanding, covering both open-source and proprietary families, and observe significant performance gaps across all tasks. Compared to human and expert performance, current MLLMs are behind—particularly in interpreting symbolic and embedded graphical elements. Our analysis further reveals key insights, including the symbolic understanding gap, the limited transferability of MLLMs to scientific domains. We release ChemTable, along with evaluation tools, to support future research in multimodal scientific understanding.
@@ 2 Related Work
Existing table recognition benchmarks mainly focus on two tasks: table structure recognition and table recognition. Early studies on table recognition relied on small but high-quality datasets such as ICDAR-2013 (ICDAR-2013). Since 2019, large-scale datasets (PubLayNet; ICDAR-2019; TableBank; PubTables-1M) have reshaped the field, enabling the deep learning era of table recognition. However, their annotations are programmatically generated and provide only table structures without cell content, limiting their utility in deeper understanding tasks. To enable deeper semantic understanding, recent datasets such as FinTabNet (fintabnet), PubTabNet (pubtabnet), and SciTSR (scitsr) incorporate logical cell locations and detailed content annotations, with TabRecSet (yang2023large) further extending coverage through multilingual and polygon-based labeling.
Beyond the general domain, scientific tables introduce richer layouts and domain-specific elements. Chemical tables, in particular, are uniquely challenging: they feature dense symbolic notation, embedded molecular structures, and implicit conventions, yet remain indispensable for reporting experimental knowledge. This complexity not only makes them valuable for practical applications but also an ideal testbed for probing the capabilities of multimodal large language models. Nevertheless, dedicated benchmarks for chemical table recognition are still lacking.
@@
Several datasets have been proposed for table understanding tasks. Among them, WikiTQ (wikitq) contains tables extracted from Wikipedia paired with natural language questions and has become a widely used early benchmark in this area. In recent years, advancements in natural language processing have extended beyond traditional homogeneous tables to incorporate additional modalities. For instance, HybridQA (hybridqa) and MMTab (MMTab) introduce multi-modal or semi-structured sources for more complex reasoning. Moreover, domain-specific benchmarks such as FinQA (finqa) (financial) and SciTab (scitab) (scientific) address unique challenges by integrating structured tables with related textual content to support complex reasoning tasks.
@@
WikiTQ wikitq
WikiSQL wikisql
HybridQA hybridqa
FinQA finqa
SciTab scitab
MMTab MMTab
@@
The ChemTable dataset is different from existing benchmarks in several key ways. It is a comprehensive, open-ended dataset focused on table recognition and understanding in chemistry, created to support scientific question-answering systems. Unlike other datasets, its questions combine visual elements, table data, and chemistry knowledge, making it necessary for models to recognize table content and structure, then reason across that information. This setup reflects real tasks that researchers face when drawing conclusions from data.
@@ 3.1.1 Table Collection
Journal Selection. To ensure both credibility and disciplinary relevance, we selected publications from top-tier chemistry journals (e.g., ACS Catalysis, JACS, Chem, Angewandte Chemie Int. Ed., and Science). The dataset covers the past decade (2015–2024), providing a balanced scope that captures recent advances while retaining representative studies across the field. All publications were systematically linked to their DOIs to guarantee traceability and to facilitate future validation or extension of the dataset. Copyright considerations regarding data usage and redistribution are detailed in Appendix S.
@@ 3.1.2 Table Annotation
For table annotation, we formed a dedicated team to identify table titles, annotations, main content, and additional captions (tablex). Each element was annotated with pixel coordinates and OCR-verified text. We also encoded the logical structure of tables, such as row-column relationships. To preserve visual meaning, we carefully recorded stylistic features like boldface, italics, and color. For table-based question answering annotation, descriptive questions are directly derived from the detailed annotations of table elements created in the previous step. Simple reasoning questions are generated using GPT-4.1 (gpt4.1) and then filtered based on difficulty. For more complex reasoning questions and visually descriptive tasks, we manually annotated 2,122 questions with the help of experienced graduate students. These questions focused on specific topics, such as reaction conditions and yields, while allowing a variety of question styles. Annotators were also encouraged to include unanswerable questions caused by missing data, vague references, or incomplete formatting. To ensure accuracy, all questions were verified by human review and model checks. Details can be found in the Appendix D.
@@ 3.2 Table Recognition
Table recognition is a fundamental step in document understanding, which aims to turn table images into structured data. This task is more challenging in chemistry because of complex layouts and specialized symbols (seeing).
@@ 3.2.2 Evaluation Protocols
Molecular Recognition: Chemical tables often include molecular structures represented as images, either embedded within cells or positioned externally. This task aims to evaluate a model’s ability to recognise and interpret such molecular graphics. The objective is to extract the corresponding SMILES (Simplified Molecular Input Line Entry System) string from a molecular diagram. This task presents unique challenges, such as fine-grained visual understanding and domain-specific symbol interpretation, which are not typically encountered in general-domain table recognition (molgrapher; generalist).
Table 2: ChemTable dataset statistics. unique tokens and QA lengths are calculated based on the Qwen2.5-7B tokenizer.
@@ 3.3.1 Task Definition
Descriptive questions aim to provide a general overview of the basic information presented in the table shown in the image. These questions include: 1) describing the main body of the table, such as its dimensions; 2) describing basic metadata of the table, including titles and notes; 3) describing domain-specific elements, such as reaction conditions in chemical reaction formulas or SMILES notation in molecular graphs; and 4) identifying certain visual features in the table, such as rows highlighted with special colors. Chemical tables often contain images, such as molecular structures, instead of plain text, making them harder to read and analyze. For example, an image in a table row can affect spacing and make it harder to recognize content. Sometimes, only certain parts of a molecule, like a red-highlighted –OH group—are shown, which adds to the difficulty for models trying to understand the table.
@@ 3.3.3 Reasoning Questions
Reasoning questions are designed to further analyze and infer information from the data presented in the table within the image. They require a comprehensive understanding to make informed judgments. These questions include: 1) Numerical and statistical reasoning, 2) Trend and change analysis, 3) Multi-hop logical reasoning, and 4) Domain-specific reasoning. Specific details are provided in the Appendix F.1. Since the data in chemical tables is often tightly linked to specific experimental conditions, molecular structures, and graphical annotations, reasoning questions not only assess the ability to comprehend explicit information but also place greater demands on domain knowledge, the integration of information across rows and columns, and the ability to reason using visual symbols.
@@ 3.3.4 Data Filtering
Diversity. We identified questions with overly repetitive structures and semantically similar phrasing to encourage a broad range of question formulations. These were rewritten using GPT-4.1 with prompt templates provided in the Appendix Q. We selected algorithms that maximised the semantic distance from the original question, enhancing the linguistic and structural diversity of the dataset.
Difficulty. We implemented a filtering strategy to ensure the dataset poses a meaningful challenge. We first conducted a single-pass QA evaluation using the Qwen-2.5-7B model (qwen2.5). For each question, we recorded whether the model was able to produce the correct answer on the first attempt. Questions that were answered correctly in one try were deemed too simple. To filter out these low-difficulty samples, we randomly discarded some of them. This approach allowed us to enrich the dataset with more challenging examples that were better suited for evaluating advanced reasoning capabilities.
@@ 4.1 Experimental Setup
To evaluate the performance of the table recognition task, we adopt the improved similarity metric based on tree edit distance (TEDS) (pubtabnet) along with the TEDS-structure indicator. Specifically, the table content may contain molecular graphs in chemical scenarios. Interpreting these molecular graphs using the simplified molecular input line entry system(SMILES) can result in structural isomorphism, where different representations or atom orders may correspond to the same molecule, making it inappropriate for TEDS to utilize normalised edit distance as a measure of cell content recognition accuracy. Therefore, for cells containing chemical molecular graphs, we replace the normalised edit distance with the Tanimoto coefficient (Tanimoto) to accurately assess the performance of table recognition. We use accuracy (ACC) as the evaluation metric for fine-grained retrieval experiments.
@@
We evaluate a diverse set of MLLMs, including open-source models InternVL3-78B (internvl3), Llama-3.2-90B (llama3.2), and Qwen2.5-VL-72B (qwen25vl), as well as proprietary models Gemini-2.5-Flash (geminiflash), GPT-4.1, GPT-4.1-mini (gpt4.1), and Claude-3.7-Sonnet (claude3). Implementation details are in Appendix L.
@@ 4.2 Experimental Results
We evaluated MLLMs on molecular formula recognition from real-world academic papers and synthetic diagrams. As shown in Figure 3, MLLMs can identify and convert molecular structures into chemical formulas, but their performance on real-world diagrams is lower than on synthetic ones, highlighting the impact of data diversity and quality on model robustness. Although MLLMs possess some chemical domain knowledge, their accuracy and reliability remain significantly inferior to specialized models (DECIMER(rajan2023decimer)). This gap underscores the need for further improvements to enable MLLMs to effectively handle domain-specific tasks in advanced scientific applications.
@@ 5.1 Experimental Setup
For tasks requiring descriptive answers, we use edit distance (edit) as the primary metric to assess recognition accuracy. For more open-ended question answering (QA) tasks, we adopt a binary evaluation strategy powered by GPT-4.1-nano (gpt4.1), which classifies each response as either correct or incorrect (mathvista; yi; alpacafarm),. The overall performance is then quantified by computing the accuracy (ACC) based on these binary classifications.
To ensure the reliability of this automated evaluation process, we randomly sampled 20% of the QA instances for manual verification. Human annotators reviewed the model outputs against reference answers and provided independent judgments. We then calculated the agreement rate between human evaluations and the binary assessments produced by GPT-4.1-nano . Implementation details are provided in the Appendix G.
@@
For table understanding, we evaluate a superset of the open-source and proprietary MLLMs introduced in Section 4.1. Concretely, we consider GPT-5, Gemini-2.5-Pro, Claude-4.5-Sonnet, GPT-4.1, Gemini-2.5-Flash, Claude-3.7-Sonnet, GPT-4.1-mini, Qwen2.5-VL-72B, InternVL3-78B, and Llama-3.2-90B. For brevity, we denote them as GPT-5, Gemini-Pro, Claude-4.5, GPT-4.1, Gemini, Claude-3.7, GPT-mini, Qwen-VL, InternVL, and Llama. We additionally include a human performance baseline (Human), consisting of human answers on the evaluation data (the collection process and evaluation protocol are detailed in Appendix L). We additionally report results for domain-specific chemistry models in Appendix K.
@@
Question Type GPT-5 Gemini-Pro Claude-4.5 GPT-4.1 Gemini Claude-3.7 GPT-mini Qwen-VL InternVL Llama Human
@@
Element
Description
Table Dimensions 74.89 74.35 76.11 73.07 73.10 75.39 68.31 71.42 70.50 70.27 -
Title Description 83.74 87.67 87.30 87.31 81.64 81.03 84.79 83.18 84.35 85.50 -
Annotation Description 93.11 89.91 87.41 92.94 81.12 68.93 81.86 90.23 73.27 81.39 -
Molecular Recognition 52.04 69.31 58.14 42.49 28.47 52.71 37.62 14.14 11.63 18.50 -
Visual
Description
Bold Description 40.53 45.81 48.93 44.27 41.22 50.38 35.88 38.93 32.82 52.73 98.99
Color Description 50.78 54.56 48.19 56.48 41.22 50.38 54.92 58.55 49.74 58.03 97.73
@@
Domain-
Specific QA
Benzene Rings Count 57.22 63.67 46.00 52.31 75.32 62.83 49.51 59.61 47.66 21.97 94.98
Yield and Conditions 90.53 92.69 90.81 89.14 90.97 89.42 89.97 85.24 74.93 82.13 93.61
Function Based 37.94 73.97 45.66 37.30 71.70 62.06 20.78 35.37 30.23 25.83 89.27
Numerical
Statistics
Value Comparison 86.44 92.00 91.85 91.80 92.00 93.60 78.40 94.40 67.14 81.45 100.00
Find Min/Max 86.18 89.62 94.79 85.85 83.18 94.39 79.44 94.39 60.94 60.95 100.00
Calculate Sum 60.65 56.43 46.32 58.33 46.43 53.68 32.63 32.63 24.07 34.12 100.00
Calculate Average 47.84 55.00 50.85 44.87 46.82 46.75 22.52 26.13 24.19 33.33 98.20
Logical
Reasoning
Trend Analysis 84.46 87.32 86.21 81.87 86.53 83.94 75.13 74.61 76.34 55.96 98.45
Multi-hop Retrieval 83.68 84.87 85.65 84.87 87.94 88.16 83.55 82.89 80.20 81.48 98.67
@@
We compare a set of representative MLLMs on the table understanding task in Table 4, covering both descriptive and reasoning questions. Below are our main findings:
(a) General reasoning strength, numerical weakness. Across general reasoning tasks such as trend analysis, value comparison, and finding min/max values, MLLMs achieve relatively high accuracy. For example, Gemini-Pro reaches 87.32 ACC on Trend Analysis, indicating that models can reliably capture basic quantitative and monotonic patterns from tables. However, performance drops sharply on arithmetic-heavy tasks: on Calculate Average, Gemini-Pro only attains 55.00 ACC, far below human performance, highlighting limitations in calculation-intensive reasoning.
(b) Descriptive tasks outperform domain-specific chemistry QA. MLLMs perform strongly on descriptive questions about table content. On Annotation Description, GPT-5 achieves 93.11 ACC, reflecting robust capabilities in recognizing and summarizing textual annotations. In contrast, accuracy decreases on domain-specific chemistry questions. Even the best model on Function Based QA, Gemini-Pro, reaches only 73.97 ACC, substantially below the human baseline, highlighting the difficulty of integrating chemical knowledge with visual table structure.
(c) Visual style interpretation remains challenging. Tasks that depend on visual or stylistic cues—such as boldface and color highlighting—remain particularly challenging. While Llama and Qwen-VL emerge as the top contenders in this category, their performance is far from ideal. Llama reaches 52.73 ACC on Bold Description, and Qwen-VL leads slightly with 58.55 on Color Description. However, compared to the near-perfect human performance (>97 ACC), this substantial gap suggests that fine-grained visual formatting is still poorly grounded in current MLLMs.
(d) Closed-source models dominate complex and domain tasks, but humans still lead. Overall, closed-source models such as GPT-5, Gemini-Pro, Claude-4.5, and GPT-4.1 dominate complex reasoning and chemistry-specific tasks. For instance, Gemini-Pro achieves 73.97 ACC on Function Based QA and over 90 ACC on Yield and Conditions, whereas most open-source models lag behind on these tasks. At the same time, strong open-source models like Qwen-VL can match or surpass proprietary ones on certain numerical subtasks (e.g., Value Comparison at 94.40 ACC). Nevertheless, humans still outperform all models on the most complex table understanding tasks, indicating a non-trivial gap to expert-level competence.